40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1493     Entry time: Fri Apr 17 11:05:22 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Thursday night locking status 
The last night, it was sort of robust to go up until arm power = 26.
The REFL_DC gain seems to change a lot around this region. So I did fine adjustments of the gain with small incremental steps of the arm power.
This work will continue.
The AutoDTT shows that the lock loss happens with an oscillation of CARM at around 100Hz. This indicates that the cross-over is the culprit.
I was also able to increase the CM UGF up to 10kHz.
ELOG V3.1.3-